PHOTOCHEMISTRY OF BIACETYL ISOLATED IN INERT GAS MATRICES

SHAWN FINNEY and \underline{C. A. BAUMANN} , Department of Chemistry, University of Scranton, Scranton, PA 18510-4626.

We report the results of ultraviolet photolysis of biacetyl ((CH_{3}CO)_{2}) trapped in solid nitrogen, oxygen, argon, krypton, and xenon. The photoproducts were characterized via infrared spectroscopy. Irradiation at 405 nm results in the production of CO, CH_{3}CO, and CH_{3}, which are the gas phase photoproducts at this wavelength ^{a} . The acetyl produced under these conditions exhibits a carbonyl stretching frequency between 1802 and 1810 cm^{−1} which is shifted from that previously observed in matrices.

aG. F. Sheats, W. A. Noyes, Jr., J. Am. Chem. Soc. 77, 1421 (1955).